A sample of 390 senior high school students applied to different state universities, wherein 117 are female. Find the 99% confidence interval of the true population proportion who applies for different state universities

1
Expert's answer
2022-06-01T13:23:27-0400

The sample proportion is computed as follows, based on the sample size "N = 390" and the number of favorable cases "X = 117"


"\hat{p}=\dfrac{X}{N}=\dfrac{117}{390}=0.3"

The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}},"

"\hat{p}+z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}})"

"=(0.3-2.5758\times \sqrt{\dfrac{0.3(1-0.3)}{390}},"

"0.3+2.5758\times \sqrt{\dfrac{0.3(1-0.3)}{390}})"

"=(0.24023, 0.35977)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.24023 < p < 0.35977," which indicates that we are 99% confident that the true population proportion "p" is contained by the interval "(0.24023, 0.35977)."